# Accretion tori around rotating neutron stars

### v1

@@
[Uncaptioned image]Monika Matuszková Accretion tori around rotating neutron starsAccretion tori around rotating neutron stars    [Uncaptioned image]Gabriel Török Accretion tori around rotating neutron starsAccretion tori around rotating neutron stars    [Uncaptioned image]Kateřina Klimovičová Accretion tori around rotating neutron starsAccretion tori around rotating neutron stars    [Uncaptioned image]Jiří Horák Accretion tori around rotating neutron starsAccretion tori around rotating neutron stars    [Uncaptioned image]Odele Straub Accretion tori around rotating neutron starsAccretion tori around rotating neutron starsAccretion tori around rotating neutron starsAccretion tori around rotating neutron stars
@@
neutron star – accretion disc – Hartle-Thorne metric
@@ 1 Introduction
Accretion disks are responsible for the strong observable radiation emitted by systems of accreting black holes (BHs) or neutron stars (NSs). Analyzing their spectral and timing properties poses an excellent opportunity to explore strong gravity (Frank et al., 2002; Kato et al., 2008) as well as super-dense matter equation of state (EoS,  Lattimer & Prakash, 2007). Here we are especially motivated by several specific challenges arising in contemporary astrophysical research in the field of X-ray variability.
@@
Throughout the history of X-ray timing observations of low-mass X-ray binaries (LMXBs) of both BH and NS types, numerous intriguing features have been revealed. The most rapid X-ray variability observed in these sources occurs at frequencies up to hundreds of Hertz, with the highest values exceeding $1.2$ kHz (Barret et al., 2006; van der Klis, 2006). Even though the rapid variability was discovered almost half a century ago, the origin of these so-called high-frequency quasi-periodic oscillations (QPOs), along with the correlated low-frequency QPOs, has never been sufficiently explained. Over the years, various models of QPOs have been proposed in an attempt to comprehend the nature of the phenomena (see, e.g., Alpar & Shaham, 1985; Lamb et al., 1985; Miller et al., 1998; Psaltis et al., 1999; Stella et al., 1999; Abramowicz & Kluźniak, 2001; Kato, 2001; Kluźniak & Abramowicz, 2001a; Wagoner et al., 2001; Titarchuk & Wood, 2002; Abramowicz et al., 2003b; Rezzolla et al., 2003b; Kluźniak et al., 2004; Zhang, 2004; Bursa, 2005; Pétri, 2005; Čadež et al., 2008; Wang et al., 2008; Mukhopadhyay, 2009; Bachetti et al., 2010; Dönmez et al., 2011; Stuchlík & Kološ, 2014; Huang et al., 2016; Le et al., 2016; Germanà, 2017; Stuchlík et al., 2020; Wang & Zhang, 2020; Smith et al., 2021; Török et al., 2022, and references therein). Most hypotheses link the QPO phenomenon to the orbital motion since the observed QPO frequencies are of the same order as those associated with the orbital motion in the innermost region of accretion disks.
@@
Numerous studies have focused on the possible relation between the QPOs and oscillatory motion of stationary pressure-supported fluid configurations (accretion tori) formed near the central compact object (e.g., Kluźniak & Abramowicz, 2001b; Kluźniak et al., 2004; Abramowicz et al., 2003a, b; Rezzolla et al., 2003a, b; Bursa, 2005; Török et al., 2005; Dönmez et al., 2011; Török et al., 2016a; de Avellar et al., 2018). In this context, Straub & Šrámková (2009) and Fragile et al. (2016) examined oscillation modes of fluid tori in the Kerr geometry, which describes spacetime around a rotating BH. In their investigation, in addition to other calculations, they evaluated the frequencies of these modes. Following their approach, we consider slowly rotating NSs for which the surrounding spacetime can be represented by the Hartle-Thorne geometry (Hartle, 1967; Hartle & Thorne, 1968). This spacetime geometry is utilized in Matuszková et al. (2024) (hereafter Paper I), where, for the first time, a comprehensive analytical description of accretion tori around rotating NSs is provided.
@@
Our paper is organized as follows. In Section 2, we briefly recall the properties of accretion tori around rotating NSs studied in Paper I. Derivation of the formulae determining these tori’s oscillation modes is the primary subject of Sections 3 and 4. From Section 5, special attention is given to investigating the behaviour of frequencies of the axisymmetric and precession modes. A complete set of formulae for calculating the frequencies of the epicyclic modes under consideration is provided within Wolfram Mathematica code and additional C++ and PYTHON libraries. In Section 6, we analyze the behaviour of these frequencies with a focus on the astrophysically relevant range of the NS parameters. We also provide simplified versions of the formulae for the frequencies of axisymmetric radial and vertical oscillations and the frequencies of the two precession modes in Section 7. Finally, in Section 8, we demonstrate the applications of our formulae on various astrophysically relevant examples, provide the reader with a short summary and state our main concluding remarks.
@@ 2 Accretion tori around rotating neutron stars
In Paper I, we describe a non-self-gravitating fluid torus orbiting around slowly rotating NS. The fluid is described by the redshift function $A\equiv u^{t}$ and its orbital velocity with respect to a distant observer which can be written as $\Omega\equiv u^{\phi}/u^{t}$. Specific angular momentum of the fluid $l$ is assumed to be constant throughout the flow ($l=l_{0}$) so that the orbital frequency of the material is Keplerian at the torus centre located at radius $r_{0}$ ($\Omega\,|_{r=r_{0}}\equiv\Omega_{0}=\Omega_{\mathrm{K}}$). Furthermore, the fluid is assumed to be polytropic with rest-mass density $\rho$, pressure $p$, total energy density $e$, and polytropic constant $n=3$.
@@
$f(r,\theta)=\frac{1}{nc_{\mathrm{s},0}^{2}}\left[\left(1+nc_{\mathrm{s},0}^{2}\right)\frac{\mathscr{E}_{0}}{\mathscr{E}}-1\right]=\mathrm{const.},$ (1)
@@
The radial and vertical torus extension is determined by the thickness parameter $\beta$, which is directly proportional to the sound speed in the torus centre $c_{\mathrm{s},0}$
@@
The NS spacetime is represented by the Hartle-Thorne geometry characterized by three parameters: NS mass $M$, dimensionless angular momentum $j$ and dimensionless quadrupole moment $q$. A detailed description of the metric, the solution for the tori, and the study of the influence of the metric parameters on the shape, size and structure of the torus are presented in Paper I. Within that paper, we also provide a Wolfram Mathematica notebook containing the full set of equations for the torus. In the following, we examine the oscillation modes of these tori.
@@ 3 Oscillations
$\left.+\frac{g^{\varphi\varphi}-\bar{\Omega}g^{t\varphi}}{1-l\Omega}\left(m-l\omega\right)^{2}\alpha f\right],$ (8)
where $\{j,k\}\in\{r,\theta\}$ are the poloidal coordinates, $\mathscr{A}\equiv A/A_{0}$, $\bar{\Omega}\equiv\Omega/\Omega_{0}$, $\bar{\omega}\equiv\omega/\Omega_{0}$, and $g$ is the determinant of the metric tensor. The normalized frequency of the oscillations $\bar{\omega}$ is the eigenvalue of the problem. The eigenfunction $W$ is related to the pressure perturbation $\delta p$ as (Abramowicz et al., 2006)
@@
where $H$ denotes the specific enthalpy. Finally, the quantity $\alpha$ in equation (8) is defined as
@@
It reflects a correction introduced by the relativistic equation of state. In all previous studies, the contribution of gas internal energy to the total mass density was neglected ($e\approx\rho$), in which case we write $\alpha=1$.
@@ 3.1 Slender torus limit
Equation (7) is a relativistic version of the Papaloizou-Pringle equation (Papaloizou & Pringle, 1984). It has no analytical solution except for the limit case of an infinitely slender torus $\left(\beta\rightarrow 0\right)$. In this limit, equation (7) becomes
@@
is the leading-order approximation of $\hat{L}$ as $\beta\rightarrow 0$. It is worth noting that equations (13) and (14) do not depend on the spacetime metric, and moreover, are identical to those in Newtonian gravity. The only quantities that depend on the gravitational field are the normalised epicyclic frequencies in the $f^{(0)}$ function in equation (3).
@@ 3.3 Non-slender tori
Blaes et al. (2006) showed that the eigenfunctions of the slender tori have polynomial dependence on $\bar{x}$ and $\bar{y}$ coordinates. The solution then can be easily found by comparing coefficients of the polynomials that appear on both sides of equation (13). The authors found all explicit expressions for the lowest order modes up to the cubic order. In particular, the epicyclic modes that are of interest here correspond to the linear eigenfunctions $W^{(0)}_{r}=\bar{x}$ and $W^{(0)}_{\theta}=\bar{y}$.
Later on, Šrámková et al. (2007) and Straub & Šrámková (2009) inspired by the earlier work of Blaes (1987) used the fact that the $\hat{L}$ operator in equation (13) is self-adjoint to derive the eigenfunctions and eigenfrequencies of the epicyclic modes in a non-slender torus case using a perturbative expansion in $\beta$ parameter. This method is also applied in this work for thicker tori in the Hartle-Thorne spacetime.
@@ 4 Perturbative expansion in torus thickness
substituting it into equation (7) and comparing the coefficients of appropriate order in $\beta$, we obtain the corresponding corrections to $W$ and $\omega$.
@@
This equation differs from equation (13) only by the presence of $\Phi^{(1)}$ on the right-hand side. Setting $\Phi^{(1)}$ to zero essentially brings us to a zeroth-order problem described by equation (13)) that has a nontrivial solution. Therefore, the right-hand side has to satisfy the solvability condition $\langle W^{(0)}\,|\,\Phi^{(1)}\rangle=0$ to assure that the solution $W^{(1)}$ exists. The solvability condition immediately provides us with the first-order correction to the eigenfrequency
@@
Similar to the tori in Kerr spacetime, the integration of the odd function over a symmetric domain here causes the scalar products in both numerators for both epicyclic modes to vanish (Straub & Šrámková, 2009). Therefore, we may write $\bar{\omega}^{(1)}=0$. The first-order correction $W^{(1)}$ can be found by solving equation (17) with $-L^{(1)}W^{(0)}$ on the right-hand side. As mentioned before, this is easily done by expanding the solution in the eigenfunctions of the zeroth-order problem and by projecting the equation on the individual eigenfunctions. We find
@@
$c_{A}=\frac{\left\langle W^{(0)}_{A}\,|\,L^{(1)}W^{(0)}\right\rangle}{2n\left[\left(\bar{\omega}^{(0)}_{A}-m\right)^{2}-\left(\bar{\omega}^{(0)}-m\right)^{2}\right]\left\langle W^{(0)}_{A}\,|\,W^{(0)}_{A}\right\rangle}.$ (21)
Since the eigenfunctions $W^{(0)}_{A}$ are polynomials, coefficients $c_{A}$ are nonzero only for a limited number of the participating modes. For the epicyclic modes, the $L^{(1)}W^{(0)}$ terms are quadratic even-parity polynomials in both $\bar{x}$ and $\bar{y}$. Consequently, the scalar product in the numerator of equation (21) does not vanish only for modes whose eigenfunctions are at most quadratic in variables $\bar{x}$ and $\bar{y}$. According to the notation introduced by Blaes et al. (2006), these are the corotation mode, two epicyclic modes, ‘X’, ‘plus’, and breathing mode. Thus, the sum in expression (20) only contains several terms.
Finally, the Papaloizou-Pringe equation (7) expanded to the second order has a similar form as the previously discussed equation (17),
@@ 5.1 Precessions
The eigenfunctions of the slender torus are assumed to be of the form $W_{r,-1}=\beta\mathcal{A}_{r}\bar{x}e^{-i\varphi}$ and $W_{\theta,-1}=\beta\mathcal{A}_{\theta}\bar{y}e^{-i\varphi}$. The eigenfrequencies are $\omega_{r,-1}=\omega_{r0}-\Omega_{0}$ and $\omega_{\theta,-1}=\omega_{\theta 0}-\Omega_{0}$. Since equations (7) and (8) are invariant under the transformation $(\omega,m)\rightarrow(-\omega,-m)$, the precessions can be equally described by $m=1$ modes with frequencies $\omega_{r,1}=\Omega_{0}-\omega_{r0}$ and $\omega_{\theta,-1}=\Omega_{0}-\omega_{\theta 0}$. In the case of thicker tori, the non-axisymmetric waves efficiently transport information between different radii so that the entire torus effectively responds as a single body with definite frequencies $\omega_{r,-1}$ and $\omega_{\theta,-1}$.

### v3

@@
[Uncaptioned image]Monika Matuszková Accretion tori around rotating neutron starsAccretion tori around rotating neutron starsAccretion tori around rotating neutron starsAccretion tori around rotating neutron stars    [Uncaptioned image]Gabriel Török Accretion tori around rotating neutron starsAccretion tori around rotating neutron stars    [Uncaptioned image]Kateřina Klimovičová Accretion tori around rotating neutron starsAccretion tori around rotating neutron stars    [Uncaptioned image]Jiří Horák Accretion tori around rotating neutron starsAccretion tori around rotating neutron stars    [Uncaptioned image]Odele Straub Accretion tori around rotating neutron starsAccretion tori around rotating neutron starsAccretion tori around rotating neutron starsAccretion tori around rotating neutron stars
@@
Stars: neutron – Accretion, accretion disks
@@ 1 Introduction
Accretion discs are responsible for the strong observable radiation emitted by systems of accreting black holes (BHs) or neutron stars (NSs). Analysing their spectral and timing properties poses an excellent opportunity to explore strong gravity (Frank et al., 2002; Kato et al., 2008) as well as super-dense matter equation of state (EoS,  Lattimer & Prakash, 2007). Here, we are especially motivated by several specific challenges arising in contemporary astrophysical research in the field of X-ray variability.
@@
Throughout the history of X-ray timing observations of low-mass X-ray binaries (LMXBs) of both BH and NS types, numerous intriguing features have been revealed. The most rapid X-ray variability observed in these sources occurs at frequencies up to hundreds of Hertz, with the highest values exceeding $1.2$ kHz (Barret et al., 2006; van der Klis, 2006). Even though the rapid variability was discovered almost half a century ago, the origin of these so-called high-frequency (HF) quasi-periodic oscillations (QPOs), along with the correlated low-frequency (LF) QPOs, has never been sufficiently explained. Over the years, various models of QPOs have been proposed in an attempt to comprehend the nature of the phenomena (see, e.g., Alpar & Shaham, 1985; Lamb et al., 1985; Miller et al., 1998; Psaltis et al., 1999; Stella et al., 1999; Abramowicz & Kluźniak, 2001; Kato, 2001; Kluźniak & Abramowicz, 2001a; Wagoner et al., 2001; Titarchuk & Wood, 2002; Abramowicz et al., 2003b; Rezzolla et al., 2003b; Kluźniak et al., 2004; Zhang, 2004; Bursa, 2005; Pétri, 2005; Čadež et al., 2008; Wang et al., 2008; Mukhopadhyay, 2009; Bachetti et al., 2010; Dönmez et al., 2011; Stuchlík & Kološ, 2014; Huang et al., 2016; Le et al., 2016; Germanà, 2017; Stuchlík et al., 2020; Wang & Zhang, 2020; Smith et al., 2021; Török et al., 2022, and references therein). Most hypotheses link the QPO phenomenon to the orbital motion since the observed QPO frequencies are of the same order as those associated with the orbital motion in the innermost region of accretion discs.
@@
Numerous studies have focused on the possible relation between the QPOs and oscillatory motion of stationary pressure-supported fluid configurations (accretion tori) formed near the central compact object (e.g., Kluźniak & Abramowicz, 2001b; Kluźniak et al., 2004; Abramowicz et al., 2003a, b; Rezzolla et al., 2003a, b; Bursa, 2005; Török et al., 2005; Dönmez et al., 2011; Török et al., 2016; de Avellar et al., 2018; Faraji & Trova, 2023). In this context, Straub & Šrámková (2009) and Fragile et al. (2016) examined oscillation modes of fluid tori in the Kerr geometry, which describes spacetime around a rotating BH. In their investigation, in addition to other calculations, they evaluated the frequencies of these modes. Following their approach, we consider slowly rotating NSs for which the surrounding spacetime can be represented by the Hartle-Thorne geometry (Hartle, 1967; Hartle & Thorne, 1968). This spacetime geometry is utilised in Matuszková et al. (2024) (hereafter Paper I), where, for the first time, a comprehensive analytical description of accretion tori around rotating NSs is provided.
@@
Our paper is organised as follows. In Sect. 2, we briefly recall the properties of accretion tori around rotating NSs studied in Paper I. Derivation of the formulae determining these tori’s oscillation modes is the primary subject of Sects. 3 and 4. From Sect. 5, special attention is given to investigating the behaviour of frequencies of the axisymmetric and precession modes. A complete set of formulae for calculating the frequencies of the epicyclic modes under consideration is provided within Wolfram Mathematica code and additional C++ and PYTHON libraries. In Sect. 6, we analyse the behaviour of these frequencies with a focus on the astrophysically relevant range of the NS parameters. We also provide simplified versions of the formulae for the frequencies of axisymmetric radial and vertical oscillations and the frequencies of the two precession modes in Sect. 7. Finally, in Sect. 8, we demonstrate the applications of our formulae on various astrophysically relevant examples, provide the reader with a short summary and state our main concluding remarks.
@@ 2 Accretion tori around rotating neutron stars
In Paper I, we describe a stationary non-self-gravitating fluid torus orbiting around a slowly rotating NS. [^1] We use units where $c=G=1$ with $c$ being the speed of light and $G$ the gravitational constant. To measure distance, we use $r_{\mathrm{G}}=GM/c^{2}$, and the ($+$)(-\,+\,+\,+)( - + + + ) metric signature is employed. Finally, the $M_{\odot}$ symbol is used for the solar mass. The fluid is described by the redshift function $A\equiv u^{t}$ and its orbital velocity with respect to a distant observer, which can be written as $\Omega\equiv u^{\varphi}/u^{t}$. Specific angular momentum of the fluid $l$ is assumed to be constant throughout the flow ($l=l_{0}$) and the specific angular momentum of the material is Keplerian at the torus centre at radius $r_{0}$, $l_{0}=l_{\mathrm{K}}(r_{0})$. Furthermore, the fluid is assumed to be polytropic with rest-mass density $\rho$, pressure $p$, total energy density $e$, and polytropic constant $n$.
[…]
@@
$f\left(r,\theta\right)=1-\frac{1}{nc_{\mathrm{s},0}^{2}}\left(1-\frac{\mathscr{E}_{0}}{\mathscr{E}}\right)=\mathrm{const.},$ (1)
@@
The radial and vertical torus proportions are determined by the thickness parameter $\beta$, which is directly proportional to the sound speed in the torus centre $c_{\mathrm{s},0}$
@@
The NS spacetime is represented by the Hartle-Thorne geometry characterised by three parameters: NS mass $M$, dimensionless angular momentum $j$ and dimensionless quadrupole moment $q$. A detailed description of the metric, the solution for the tori, and the study of the influence of the metric parameters on the shape, size and structure of the torus are presented in Paper I. Within that paper, we also provide a Wolfram Mathematica notebook containing the full set of equations for the torus. In the following, we examine the oscillation modes of these tori.
@@ 3 Oscillations
$\left.-\frac{g^{\varphi\varphi}-\Omega g^{t\varphi}}{1-l\Omega}\left(m-l\omega\right)^{2}\alpha f\right],$ (8)
where $\{j,k\}\in\{r,\theta\}$ are the poloidal coordinates, $\mathscr{A}\equiv A/A_{0}$, $\bar{\Omega}\equiv\Omega/\Omega_{0}$, $\bar{\omega}\equiv\omega/\Omega_{0}$, and $g$ is the determinant of the metric tensor. The normalised frequency of the oscillations $\bar{\omega}$ is the eigenvalue of the problem. The eigenfunction $W$ is related to the pressure perturbation $\delta p$ as (Abramowicz et al., 2006)
@@
where $H$ denotes the specific enthalpy. Finally, the quantity $\alpha$ in Eq. (8) reflects a correction introduced by the relativistic equation of state. In all previous studies, the contribution of gas internal energy to the total mass density was neglected ($e\approx\rho$), corresponding to $\alpha=1$. Here, we use a consistent expression
@@ 3.1 Slender torus limit
Equation (7) is a relativistic version of the Papaloizou-Pringle equation (Papaloizou & Pringle, 1984). It has no analytical solution except for the limit case of an infinitely slender torus $\left(\beta\rightarrow 0\right)$. In this limit, Eq. (7) becomes
@@
is the leading-order approximation of $\hat{L}$ as $\beta\rightarrow 0$. It is worth noting that Eqs. (13) and (14) do not depend on the spacetime metric and, moreover, are identical to those in Newtonian gravity. The only quantities that depend on the gravitational field are the normalised epicyclic frequencies in the $f^{(0)}$ function in Eq. (3).
@@ 3.3 Non-slender tori
Blaes et al. (2006) showed that the eigenfunctions of the slender tori have polynomial dependence on $\bar{x}$ and $\bar{y}$ coordinates. The solution then can be easily found by comparing coefficients of the polynomials that appear on both sides of Eq. (13). The authors found all explicit expressions for the lowest order modes up to the cubic order. In particular, the epicyclic modes that are of interest here correspond to the linear eigenfunctions $W^{(0)}_{r}=\bar{x}$ and $W^{(0)}_{\theta}=\bar{y}$.
Later on, Šrámková et al. (2007) and Straub & Šrámková (2009) inspired by the earlier work of Blaes (1987) used the fact that the $\hat{L}$ operator in Eq. (13) is self-adjoint to derive the eigenfunctions and eigenfrequencies of the epicyclic modes in a non-slender torus case using a perturbative expansion in $\beta$ parameter. This method is also applied in this work for thicker tori in the Hartle-Thorne spacetime.
@@ 4 Perturbative expansion in torus thickness
substituting it into Eq. (7) and comparing the coefficients of appropriate order in $\beta$, we obtain the corresponding corrections to $W$ and $\omega$.
@@
This equation differs from Eq. (13) only by the presence of $\Phi^{(1)}$ on the right-hand side. Setting $\Phi^{(1)}$ to zero essentially brings us to a zeroth-order problem described by Eq. (13)) that has a nontrivial solution. Therefore, the right-hand side has to satisfy the solvability condition $\langle W^{(0)}\,|\,\Phi^{(1)}\rangle=0$ to assure that the solution $W^{(1)}$ exists. The solvability condition immediately provides us with the first-order correction to the eigenfrequency
@@
Similar to the tori in Kerr spacetime, the integration of the odd function over a symmetric domain here causes the scalar products in both numerators for both epicyclic modes to vanish (Straub & Šrámková, 2009). Therefore, we may write $\bar{\omega}^{(1)}=0$. The first-order correction $W^{(1)}$ can be found by solving Eq. (17) with $-L^{(1)}W^{(0)}+4nm\left(\bar{\omega}^{(0)}-m\right)\bar{\Omega}^{(1)}W^{(0)}$ on the right-hand side. As mentioned before, this is easily done by expanding the solution in the eigenfunctions of the zeroth-order problem and by projecting the equation on the individual eigenfunctions. We find
@@
$c_{A}=\frac{\left\langle W^{(0)}_{A}\,|\,L^{(1)}W^{(0)}\right\rangle-4nm\left(\bar{\omega}^{(0)}-m\right)\left\langle W^{(0)}_{A}\,|\,\bar{\Omega}^{(1)}W^{(0)}\right\rangle}{2n\left[\left(\bar{\omega}^{(0)}_{A}-m\right)^{2}-\left(\bar{\omega}^{(0)}-m\right)^{2}\right]\left\langle W^{(0)}_{A}\,|\,W^{(0)}_{A}\right\rangle}.$ (21)
Since the eigenfunctions $W^{(0)}_{A}$ are polynomials, coefficients $c_{A}$ are nonzero only for a limited number of the participating modes. For the epicyclic modes, the $L^{(1)}W^{(0)}$ terms are quadratic even-parity polynomials in both $\bar{x}$ and $\bar{y}$. Consequently, the scalar product in the numerator of Eq. (21) does not vanish only for modes whose eigenfunctions are at most quadratic in variables $\bar{x}$ and $\bar{y}$. According to the notation introduced by Blaes et al. (2006), these are the corotation mode, two epicyclic modes, ‘X’, ‘plus’, and breathing mode. Thus, the sum in expression (20) only contains several terms.
Finally, the Papaloizou-Pringe Eq. (7) expanded to the second order has a similar form as the previously discussed Eq. (17),
@@ 5.1 Precessions
The eigenfunctions of the slender torus are assumed to be of the form $W_{r,-1}=\beta\mathcal{A}_{r}\bar{x}e^{-i\varphi}$ and $W_{\theta,-1}=\beta\mathcal{A}_{\theta}\bar{y}e^{-i\varphi}$. The eigenfrequencies are $\omega_{r,-1}=\omega_{r0}-\Omega_{0}$ and $\omega_{\theta,-1}=\omega_{\theta 0}-\Omega_{0}$. Since Eqs. (7) and (8) are invariant under the transformation $(\omega,m)\rightarrow(-\omega,-m)$, the precessions can be equally described by $m=1$ modes with frequencies $\omega_{r,1}=\Omega_{0}-\omega_{r0}$ and $\omega_{\theta,-1}=\Omega_{0}-\omega_{\theta 0}$. In the case of thicker tori, the non-axisymmetric waves efficiently transport information between different radii so that the entire torus effectively responds as a single body with definite frequencies $\omega_{r,-1}$ and $\omega_{\theta,-1}$.
Figure 1: The axisymmetric oscillation frequencies calculated for different $\beta$ and NSs with $j=0.2$, and two particular values of $q/j^{2}$, $q/j^{2}=1.5$ (extreme NS compactness), and $q/j^{2}=10$ (large NS oblateness). Left: The radial mode. Right: The vertical mode. The curves indicating particular values of $\beta$ are normalised to the maximal value $\beta=\beta_{\mathrm{max}}$. The critical radius limiting the existence of cusp tori is denoted by the dotted vertical line marked as $r_{\mathrm{mc}}$. The maximal value $\beta=\beta_{\mathrm{max}}$ matches the cusp value when $r<r_{\mathrm{mc}}$, and the value implying tori extending to infinity when $r>r_{\mathrm{mc}}$. For the absolute value of $\beta$, only one specific curve is drawn in each panel, marked in grey, $\beta=\sqrt{0.15}$, which makes it possible to distinguish frequencies that correspond to relatively large tori, as they are displayed below this curve. The dotted vertical lines in each panel marked as $r_{mc}$ denote the radii limiting the existence of cusp tori – any torus with a centre above this radii does not exhibit the cusp. An analogous notation is later used in several other figures.